### v2

@@ Appendix G Appendix: Five-Digit Addition Experiments
Appendix H Appendix: Addition Interchange Interventions
To test the hypothesis 3 mapping of the mathematical framework (casual abstraction) to the model attention heads, various “interchange interventions” (aka activation patching) experiments were performed on the model, where
@@ Appendix H Appendix: Addition Interchange Interventions
Using this approach we obtained the findiings in Tab. 12.
Table 12: Interchange Interventions experiments used activation patching to test the claims addition hypothesis 3 made for each attention head. Experimental results are consistent with hypothesis 3 for all nodes.
@@
P8.L0.H1 and MLP
D2.C = TriCase( D2, D’2 ) impacting A4 and A5 accuracy
Based on D2 and D’2. Triggers on a D2 carry value. Provides carry bit used in A5 and A4 calculation.
P9.L0.H1 and MLP
D1.C = TriCase( D1, D’1 ) impacting A5, A4 and A3 accuracy
Based on D1 and D’1. Triggers on a D1 carry value. Provides carry bit used in A5, A4 and A3 calculation.
P10.L0.H1 and MLP
D1.C2 = TriAdd( D1.C, TriCase(D0, D’0) ) impacting A5, A4, A3 and A2 accuracy
Based on D0 and D’0. Triggers on a D0 carry value. Provides carry bit used in A5, A4, A3 and A2 calculation.
P11.L0.H1 and MLP
D3.C4 = TriAdd( TriCase( D3, D’3 ), TriAdd( D2.C, D1.C2)) impacting A5 accuracy
Based on D3 and D’3. Triggers on a D3 carry value. Provides carry bit used in A5 calculations.
P11.L0.H2 and MLP
D4.C = TriCase(D4, D’4) impacting A5 accuracy A4
Based on D4 and D’4. Triggers on a D4 carry value. Provides carry bit used in A5 calculation.
P12.L0.H0+H2 and MLP
D4.BA = (D4 + D’4) % 10 impacting A4 accuracy A4
Sums D4 and D’4. Impacts A4.
P13.L0.H0+H2 and MLP
D3.BA = (D3 + D’3) % 10 impacting A3 accuracy
Sums D3 and D’3. Impacts A3.
P14.L0.H0+H2 and MLP
D2.BA = (D2 + D’2) % 10 impacting A2 accuracy
Sums D2 and D’2. Impacts A2.
P14.L0.H1 and MLP
(D1 + D’1) / 10 + P10.D1.C2 info impacting A2 accuracy
Calculates P10.D1.C1 but add P10.D1.C2 info when D1.C != D1.C2. Impacts A2
P15.L0.H0+H2 and MLP
D1.BA = (D1 + D’1) % 10 impacting A1 accuracy
Sums D1 and D’1. Impacts A1.
P15.L0.H1 and MLP
D0.MC = (D0 + D’0) / 10 impacting A1 accuracy
Triggers when D0 + D’0 >10. Impacts A1 digit by 1
P16.L0.H0+H2 and MLP
D0.BA = (D0 + D’0) % 10 impacting A0 accuracy
Sums D0 and D’0. Impacts A0.
Appendix I Appendix: Six-Digit Addition
We trained a very low loss 6-digit Addition model. We compared the 5-digit and 6-digit model algorithms. Figures 6 and  8 are very instructive. So are Figures 7 and  17. We concluded that the 6-digit model used the same logical algorithm as the 5-digit model. There were some implementation differences:
@@ Appendix I Appendix: Six-Digit Addition
  • The 5-digit model used 2 attention heads per digit to do the BA calculation, whereas 6-digit model only uses one (and so is more compact).
@@
  • The PCA trigrams have difference appearances in the two models. Figures 13 and  18.
  • For 6-digits, for A2 (only), the BA calculation is done in P18, but the MC calculation is not in P18. This is unique. The 6-digit model has optimised out the MC and instead relies solely on the accurate D1.C2 value (and so is more compact).
Figure 16: This map shows the input tokens each attention head attends to at each token position. The model predicts the first answer digit A6 in token position P14, and all digit-pair sums (e.g. D2 D’2) are attended to in P11 to P14. Note the similarity to Figure 5.
Figure 17: This map shows the answer digit(s) A0 .. A5 impacted when we ablate each useful node in the 6-digit 2-layer 3-head addition model. Note the similarity to Figure 7.
Figure 18: For 9 attention heads, for exactly 1 answer digit $A_{n}$each, PCA shows these interpretable results for the 6-digit Addition model. The 3 dot colours show the TriCase value of each question. The PCA data and TriCase quanta are both tri-state and strongly correlated.
Appendix J Appendix: Six-Digit Subtraction
Integer subtraction (e.g. 654321-789012=-0134691) and integer addition have similarities. Subtraction is conceptually more complex than integer addition. While $a+b$ and $b+a$ give the same answer, generally $a-b$ and $b-a$ give difference answers. In human subtraction the algorithm used to calculate $a-b$ changes depending on whether $a>b$ or $a<b$.
In our transformer-based models, Subtraction and Addition have some similar quanta. For example:
  • In addition, the first answer digit is always and “1” or “0” and this token is the hardest to calculate accurately because of the “cascading carry 1” edge case. In subtraction, the first answer token is always “+” or “-” and this token is the hardest to calculate accurately because of the “cascading borrow 1” edge case.
  • If $a>b$ then the algorithm to calculate $a-b$ and $a+b$ have strong structural similarities, allowing per-digit calculations with one sub-task moving data between digits (“carry 1” or “borrow 1”)
  • If $a<b$ then the algorithm to calculate $a-b$ can use the mathematical identity that $1a-b=-(b-a)$. Our intuition is that the model will calculate if $a<b$, and if so will reuse the $a>b$ circuit, then negate the answer.
We created a very-low-loss 6-digit addition model. Figure 8 shows the lowest quanta impacted when ablating each node. We created a “mixed operations” model (6-digit, 3-layer, 4-head) to do both Addition and Subtraction. We initialised the untrained model with the addition model and trained it (with 80% subtraction and 20% addition questions) to be very low loss. This mixed model can predict one million addition and one million subtraction questions without error.
Figure 19: The per-digit log-loss curves for a 6-digit 3-layer 4-head mixed model trained to perform subtraction and addition after being initialised with a very low loss addition model.
To investigate what classes of questions the model is always getting wrong, we expanded our existing addition-specific metrics, to include subtraction-specific question categorisations (see Sec. 4.3. Figures 9 and  10 maps these quanta against nodes showing which are used for addition and subtraction.
@@
The model contains a new sub-task that stands out: The algorithm relies on calculations done at token position P0, when the model has only seen one question token! What information can the model gather from just the first token? Intuitively, if the first token is a “8” or “9” then the first answer token is more likely to be a “+” (and not a “-”). The model uses this heuristic even though this probabilistic information is sometimes incorrect and so will work against the model achieving very low loss.
Figure 20: This map shows the answer digit(s) A0 .. A7 impacted when we ablate each useful node in the 6-digit 3-layer 4-head mixed model. A7 is the sign of the answer and calculated first. Note the similarity to the 5 and 6-addition Figures 7 and  17.

### v7

@@ Appendix I Appendix: Addition Hypothesis 2
Appendix K Appendix: Addition Interchange Interventions
To test the hypothesis 3 mapping of the mathematical framework (casual abstraction) to the model attention heads, various “interchange interventions" (aka activation patching) experiments were performed on the model, where
@@
Using this approach we obtained the findings in Tab. 12.
@@
P8.L0.H1 and MLP A2.ST = TriCase( D2, D’2 ) impacting A4 and A5 accuracy Based on D2 and D’2. Triggers on a A2 carry value. Provides carry bit used in A5 and A4 calculation.
P9.L0.H1 and MLP A1.ST = TriCase( D1, D’1 ) impacting A5, A4 and A3 accuracy Based on D1 and D’1. Triggers on a A1 carry value. Provides carry bit used in A5, A4 and A3 calculation.
P10.L0.H1 and MLP A1.ST2 = TriAdd( A1.ST, TriCase(D0, D’0) ) impacting A5, A4, A3 and A2 accuracy Based on D0 and D’0. Triggers on a A0 carry value. Provides carry bit used in A5, A4, A3 and A2 calculation.
P11.L0.H1 and MLP A3.ST4 = TriAdd( TriCase( D3, D’3 ), TriAdd( A2.ST, A1.ST2)) impacting A5 accuracy Based on D3 and D’3. Triggers on a A3 carry value. Provides carry bit used in A5 calculations.
P11.L0.H2 and MLP A4.ST = TriCase(D4, D’4) impacting A5 accuracy A4 Based on D4 and D’4. Triggers on a A4 carry value. Provides carry bit used in A5 calculation.
P12.L0.H0+H2 and MLP A4.SA = (D4 + D’4) % 10 impacting A4 accuracy A4 Sums D4 and D’4. Impacts A4.
P13.L0.H0+H2 and MLP A3.SA = (D3 + D’3) % 10 impacting A3 accuracy Sums D3 and D’3. Impacts A3.
P14.L0.H0+H2 and MLP A2.SA = (D2 + D’2) % 10 impacting A2 accuracy Sums D2 and D’2. Impacts A2.
P14.L0.H1 and MLP (D1 + D’1) / 10 + P10.A1.ST2 info impacting A2 accuracy Calculates P10.A1.ST1 but add P10.A1.ST2 info when A1.ST != A1.ST2. Impacts A2
P15.L0.H0+H2 and MLP A1.SA = (D1 + D’1) % 10 impacting A1 accuracy Sums D1 and D’1. Impacts A1.
P15.L0.H1 and MLP A0.SC = (D0 + D’0) / 10 impacting A1 accuracy Triggers when D0 + D’0 >10. Impacts A1 digit by 1
P16.L0.H0+H2 and MLP A0.SA = (D0 + D’0) % 10 impacting A0 accuracy Sums D0 and D’0. Impacts A0.
Table 12: Interchange Interventions experiments used activation patching to test the claims addition hypothesis 3 made for each attention head in a sample mixed model. Experimental results are consistent with hypothesis 3 for all nodes.
+ve Sub Contains -ve Sub Contains Like
M0 MD N/A N/A S0
M1 MD,MB N1 ND,NB S1
M2 MD,MBx2 N2 ND,NBx2 S2
M3 MD,MBx3 N3 ND,NBx3 S3
M4 MD,MBx4 N4 ND,NBx4 S4
Table 13: We define “positive-answer subtraction" and “negative-answer subtraction" calculation complexity quanta that parallel the addition quanta.
Appendix L Appendix: N-Digit Addition
The addition models perform addition accurately. Visualizations that provided insights into the behavior of the model, aiding our interpretation of the algorithm, are below:
Figure 10: This map shows the simpliest (lowest complexity) quanta S0, S1, etc impacted when we ablate each node in the 5-digit 2-layer 3-head addition model. To answer S0 questions, only the S0 nodes are used. To answer S1 questions, S0 and S1 nodes are used, etc.
Some notes about the models:
@@
  • Some models use 2 attention heads per digit to do the SA calculation, whereas some models only uses one (and so are more compact).
@@
  • The PCA trigrams have difference appearances in different models (but the same interpretable clusters). Refer Figures 8
  • Per answer digit, some models use the SC calculation, whereas some models optimize it out and rely solely on the $ST$ value (and so are more compact).
Appendix M Appendix: Mixed Model Initialization
We experimented with three approaches to re-using the trained addition model in the “mixed" (addition and subtraction) model:
  • Initialize Only: Initialize the untrained mixed model with the addition model weights before training begins.
  • Freeze Attention: As per “Initialize Only", but also every 100 training steps recopy the attention head weights from the addition model into the partially-trained mixed model.
  • Freeze All: As per “Initialize Only", but also every 100 training steps recopy the entire addition model (attention heads and MLP layers) into the partially-trained mixed model.
Our intuition was that “Initialize Only" would give the mixed model the most freedom to learn new algorithms, but that the “Freeze Attention" and “Freeze All" approaches would make the resulting trained mixed model easier to interpret (as we could reuse our addition model insights).
After experimentation we found that the “Initialize Only" approach was the only one that quickly trained to be able to do both addition and subtraction accurately. We concluded that the other two methods constrainws the model’s ability to learn new algorithms too much.
We also experimented with “where" in the model we inserted the addition (6-digit, 2-layer, 3-head) model into the slightly larger (6-digit, 3-layer, 4-head) mixed model. That is, do we initialize the first 2 layers or the last 2 layers of the mixed model? Also do we initialize the first 3 attention heads or the last 3 attention heads of the mixed model? Our intuition was that initializing the first layers and heads would be more likely to cause the model to re-use the addition circuits adding interpretability, so we used this approach.
Appendix N Appendix: N-Digit Subtraction
The mixed models perform addition and subtraction accurately. Visualizations that provided insights into the behavior of the model, aiding our interpretation of the algorithm, are below:
Figure 12: This map of a sample 6-digit mixed model shows the 98 nodes used to predict answers to addition (S), positive-answer subtraction (M) and/or negative-answer subtraction (N) questions. Before training the mixed model, 48 nodes were initalized pre-training with a smaller addition model’s weights. These are have a red border. During mixed model training, 39 of 48 of the initalized monosemantic nodes were generalized (become poly-semantic) and now help predict two or three question classes.
Figure 14: This map shows the simpliest complexity quanta M0, M1, etc used in each useful node of the 6-digit 3-layer 4-head mixed model for subtraction questions with positive answers.
Some notes about the mixed models:
  • All the notes about the addition model (above) also apply to the mixed model.
@@
  • The model contains a new sub-task that stands out: The algorithm relies on calculations done at token position P0, when the model has only seen one question token! What information can the model gather from just the first token? Intuitively, if the first token is a “8" or “9" then the first answer token is more likely to be a “+" (and not a “-"). The model uses this heuristic even though this probabilistic information is sometimes incorrect and so will work against the model achieving very low loss.